OpenStudy (anonymous):

one number is 5 more than another number. if you add 8 to 2 times the first number, the result is 3 times the second number. (choose the system of equations in x and y describing the situations) a. x+y=5 2x-3y=-8 B. x-y=5 -2x+x-y=5 C.x-y=5 2x-3y=-8 D. x-y=5 2x-3y=8

OpenStudy (anonymous):

one number is 5 more than another number. this gives x-y = 5

OpenStudy (anonymous):

so that elemanates A

OpenStudy (anonymous):

if you add 8 to 2 times the first number, the result is 3 times the second number this can be written as 2x + 8 = 3y or 2x - 3y = -8

OpenStudy (anonymous):

so ans is C.x-y=5 2x-3y=-8

OpenStudy (anonymous):

now I understand than you so much

OpenStudy (anonymous):

yes it does eliminate A.